OpenStudy (anonymous):

4|4-7v|+8=108 solve the equation so there are two answers for v. I got the first v, but when the equation is switched to -4|4-7v|+8=108 the answer didn't for correctly back in when I checked it. Help please?

OpenStudy (a_clan):

4|4-7v|+8=108. This equation can have two forms: 1)....4(4-7v)+8=108 2)....-4(4-7v)+8=108 Solve the two equations. These will give correct answers, if your calculations are right.

OpenStudy (anonymous):

So is the second answer 29/7